OpenStudy (anonymous):

4b/b-2 -[subtract] 1/3= 13/3b-6

OpenStudy (gw2011):

If the problem is as follows: 4b/(b-2) - 1/3 = 13/(3b-6) then the solution is: 4b/(b-2) - 13/(3b-6) = 1/3 4b/(b-2) - 13/3(b-2) = 1/3 Multiply both sides by 3 to get: (4b)(3)/(b-2) - (13)(3)/3(b-2) = (1/3)(3) 12b/(b-2) - 39/3(b-2) = 1 Multiply both sides by 3(b-2) and you get: (12b)(3) - 39 = (1)(3)(b-2) 36b - 39 = 3b - 6 36b - 3b = 39 - 6 33b = 33 b = 1